OpenStudy (anonymous):

find an equation for a line that goes through (1,8) and (-1,-2)

OpenStudy (anonymous):

first you find the slope

OpenStudy (anonymous):

once you find the slop use this formula and put your slope where th m is on this formula y = mx + b

OpenStudy (anonymous):

now take either (1, 8) or (-1, -2) and plug in the values of x and y

OpenStudy (anonymous):

remember the answer you get cause you will still need it

OpenStudy (anonymous):

you should be able to figure the rest out

OpenStudy (anonymous):

ill provide an example i found that is easy to understand

OpenStudy (anonymous):

ok so your to points for example are (5,13) and (8,19) first you find the slope: (19 - 13) / (8 - 5) = 6 / 3 = 2 SLOPE = 2, remember that. now write down: y = mx + b use the slope we found: y = 2x + b ***now take either (5, 13) or (8, 19) and plug in the values of x and y*** we'll use (5, 13): y = 2x + b 13 = 2(5) + b 13 = 10 + b b = 3 we have now found: slope = 2 and b = 3 so we write down y = mx + b again, and plug in for the info we found. y = mx + b y = 2x + 3 <--- answer

OpenStudy (anonymous):

use that example to figure out your answer